# Towards General Industrial Intelligence: A Survey on Industrial IoT-Enhanced Continual Large Models

### v1

@@ III-B LMs as a Tuning Predictor
Figure 16: Different types of generative models.
Several fine-tuning strategies have been developed to meet these needs. Here, we focus on two primary adaptation methods: alignment fine-tuning and efficient fine-tuning. Alignment fine-tuning is designed to adjust the model to better match the semantics and operational standards of specific tasks, while Efficient fine-tuning aims to reduce the computational resources and time required while maintaining model performance. Each method has its own application scenarios and advantages, providing flexible adjustment tools for the practical deployment of LMs in industrial settings.
III-B1 Alignment Fine-tuning
In this section, we introduce Alignment fine-tuning, a fine-tuning approach that enhances the alignment of pre-trained LMs with specific task instructions. Alignment fine-tuning primarily consists of three sub-strategies: Instruction Tuning, Model Refinement, and Model Alignment.
$\bullet$ Instruction Tuning. Instruction Tuning (IT) is a crucial technique for adapting pre-trained LLMs to specific tasks and user preferences in various domains, ensuring alignment with nuanced requirements.
@@ III-B1 Alignment Fine-tuning
$\bullet$ Model Refinement. Like humans, LLMs are prone to errors, such as producing inaccurate translations or retaining outdated information. Ensuring that LLMs remain accurate and up-to-date typically involves direct fine-tuning of the model. While effective, this method can be time-consuming and risk degrading performance on tasks the model has previously learned. To overcome these challenges, the concept of Model Refinement (MR) [99] has been developed as a technique to correct specific errors without disrupting the primary functionalities of the model. This approach, initially guided by a “reliability-locality-efficiency” principle, implements an efficient editing process using gradient descent. Subsequent research [100, 101] have expanded on this foundation, applying gradient decomposition techniques to larger models such as GPT-J-6B and T5-XXL. These methods focus on updating a subset of model parameters to modify the labels of specific inputs and integrate editing through retrieval mechanisms.
@@
Figure 18: An example of IT data.
$\bullet$ Model Alignment. Model Alignment (MA) is a fundamental concept in the deployment of AI systems, aimed at ensuring that the actions and outputs of AI models are in harmony with human values, ethics, and preferences. This process involves the use of mathematical models, algorithmic adjustments, and iterative feedback to refine AI behaviours, enabling the model to adapt continuously to a stream of incoming data and remain up-to-date.
@@
III-B2 Efficient Fine-tuning
@@
Density Efficient. With increasing model sizes, traditional fine-tuning of all parameters becomes increasingly challenging. PEFT methods have gained popularity for effectively adapting pre-trained LMs. Recent research has identified the presence of activation sparsity within the Multilayer Perceptron (MLP) blocks of transformers. Lower activation density enables efficient model inference on hardware optimized for sparsity. Building on this insight, a new method called DEFT [107] has been proposed to reduce loss by promoting higher activation sparsity (or lower activation density) in pre-trained models.
Memory Efficient. Traditional fine-tuning methods rely on backpropagation algorithms. However, as the model size increases, they demand exponentially more memory. Zeroth-order methods theoretically estimate gradients with only two forward passes but are typically very slow when optimizing LMs. MeZO [108] enhances the traditional Zeroth-order Stochastic Gradient Descent method by adapting it to in-place operations, maintaining the memory usage for fine-tuning LMs at the same level as that during inference.
Parameter Efficient. Parameter-efficient fine-tuning methods adapt a large pre-trained LM to specific tasks by fine-tuning only a small subset of parameters while leaving most of the model’s parameters fixed. Below are several common strategies used for parameter-efficient fine-tuning:
@@
Discussion. This section (including Section III-B1 and Section III-B2) explores the application of LMs as tuning predictors in IIoT-driven general industrial intelligence. By introducing two major strategies—alignment fine-tuning and efficient fine-tuning—it provides an in-depth analysis of methods such as instruction tuning, model refinement, and model alignment, examining how they enhance the adaptability of LMs in various industrial scenarios.
However, challenges in industrial applications remain, particularly in addressing catastrophic forgetting and the high costs associated with training. Future research will need to incorporate advanced optimization strategies, such as in-context learning and chain-of-thought [116] reasoning, to overcome these limitations and ensure stability and reliability in predictions. Moreover, as edge computing and communication technologies continue to advance, the ability to continuously optimize LMs’ performance in dynamic and multimodal data environments will be crucial for the further development of general industrial intelligence.
III-C LMs as a Non-Tuning Predictor
@@
Figure 21: In-context Learning for classification tasks. Several demonstrations containing both sources (comments for a book) and target classes (sentiments for the book) are fed into the LLM with the test input/query that only includes comment. The LLM can learn from the demonstrations and predicts the result without updating its parameters.
Prompt-based Methods. A notable method translates natural language instructions into executable robot actions [82]. This approach proposes customizable input prompts that seamlessly integrate into robotic execution systems or visual recognition programs, facilitating multi-step task plans while mitigating token limitation constraints of ChatGPT. In autonomous driving, LimSim++ [117] provides diverse types and modalities of prompt inputs, extracting road network and vehicle surroundings information at each decision frame, which is then conveyed in natural language to the driving agent (e.g., GPT-4).
Tokenizer-based Methods. The introduction of LiDAR-LLM [118] leverages LLMs’ reasoning capabilities to comprehend outdoor 3D scenes fully. Addressing the scarcity of paired 3D LiDAR and text data, a three-stage training strategy generates datasets that align the 3D modality with the language embedding space of LLMs. Initially, LLM tokenizers were not designed to handle numerical data or recognize associations in time series effectively. This challenge was addressed through lightweight embedding layers and refined prompt engineering [119]. LLMTime [120] utilizes innovative tokenization techniques, transforming tokens into flexible, continuous-valued representations, enabling unadjusted LLMs to achieve or surpass traditional performance in zero-shot prediction tasks.
@@
In-Context Learning (ICL). Introduced in GPT-3 [121], ICL, initially popularized by GPT-3, enables LLMs to make accurate predictions using minimal demonstrations, bypassing the need for gradient updates. This update-free feature not only lowers adaptation costs but also broadens the model’s applicability across various tasks. The efficacy of ICL hinges on both the pre-training of the model on extensive datasets and the careful selection of demonstrations for task-specific predictions. Studies have identified several critical factors influencing the performance of ICL, including the label space, input text distribution, and the format of demonstrations [122]. Moreover, structural similarity, diversity, and complexity within demonstrations are also significant [123]. Recent research by Coda-Forno et al. [124] has shown that the capabilities of LLMs in ICL can be recursively enhanced through the ICL process itself. Fig. 21 shows an example of in-context learning.
Discussion. This section explores the application of closed-source large models in industrial environments through structured data preprocessing and model interaction in non-tuning predictor scenarios. As edge computing and communication technologies evolve, effectively leveraging non-tuning capabilities of large models to handle multimodal data and real-time tasks in industrial settings will be critical for optimizing general industrial intelligence systems. Future research will focus on refining prompt engineering and improving tokenization techniques to enhance the adaptability and stability of non-tuning large models in complex industrial scenarios.
IV Large Models on General Industrial Intelligence
In the domain of LMs, traditional pre-trained often relies on processing massive datasets, such as corpora comprising billions of words. Faced with new data, these models typically require retraining from scratch. However, distribution shifts introduced by new data often degrade performance on previous tasks or impede adaptation to new data. Employing CL strategies for large models can significantly reduce computational resource consumption and enhance efficiency.
The essence of CL strategies is to enable large models to adapt to data evolving over time and across domains, thereby maintaining the accuracy and freshness of their knowledge. This strategy provides a robust framework for large models, enhancing their adaptability to real-time changes in industrial environments. Challenges in CL include:
$\bullet$ Catastrophic Forgetting: Models frequently forget previously learned tasks when adapting to new ones, particularly evident in horizontal CL as models switch between tasks with varying data distributions.
$\bullet$ Extended Task Sequences: Facing time-series data or evolving task structures, models require continuous updates, which may lead to forgetting earlier learned tasks.
$\bullet$ Distributional Shifts: Sudden changes in data distribution pose greater demands on the model’s generalization capabilities and adaptability.
$\bullet$ Unknown Task Learning: Incorporating new and previously unseen tasks requires robust mechanisms to adapt without forgetting existing tasks.
$\bullet$ Incorporation of New Modalities: As industrial environments integrate new data modalities (e.g., visual, audio), models need the ability to align and learn from these modalities without disrupting existing knowledge.
Recent research have shown that with simple and scalable CL strategies, large models can be effectively updated using only a fraction of the computational resources, achieving performance comparable to that of full retraining [125]. While these advancements are promising, further exploration is needed to assess their effectiveness, particularly in the context of LLMs applied to industrial environments. This section reviews classic CL methods, investigates their application in pre-trained models, and explores the latest online adaptation techniques relevant to general industrial intelligence.
IV-A Classical Continual Learning
CL aims to enable models to adapt to new information continuously while retaining mastery over previously acquired knowledge, as illustrated in Fig. 22. The objectives of CL include reducing the forgetting of previous knowledge and leveraging prior knowledge to expedite new task learning, thereby improving efficiency and reducing computational resource consumption.
CL is currently used in various industrial domains, such as fault diagnosis [126] and target detection [127]. Empirical studies such as those by Luo et al. [128] highlight the challenges of catastrophic forgetting in LLMs and suggest methods to mitigate these effects through innovative CL strategies. These include regularization methods that aim to preserve previous knowledge by adding constraints during the learning of new tasks, and replay methods that involve reusing old data to reinforce memory. Moreover, parameter isolation techniques provide another avenue for CL by dedicating specific parameters to particular tasks, thus protecting learned knowledge from being overwritten.
In summary, balancing learning plasticity and memory stability is crucial in CL, requiring strategies that can adapt models to new data without losing the ability to perform well on previously learned tasks. As shown in Fig. 23, we mainly discuss classical CL as (1) Regularization-based methods, (2) Replay-based methods, and (3) Parameter Isolation methods.

### v2

@@ IV-B LMs as the Tuning Predictor
IV-B1 The Goal of Fine-tuning
Fine-tuning LLMs serves multiple objectives, including adapting models to specific tasks, correcting errors, and ensuring alignment with human values. These objectives can be categorized into three key areas: Instruction Tuning, Model Refinement, and Model Alignment, each addressing distinct challenges and advancements in fine-tuning methodologies.
Instruction Tuning. Instruction Tuning (IT) is a crucial technique for adapting pre-trained LLMs to specific tasks and user preferences in various domains, ensuring alignment with nuanced requirements.
@@ IV-B1 The Goal of Fine-tuning
Model Refinement. Like humans, LLMs are prone to errors, such as producing inaccurate translations or retaining outdated information. Ensuring that LLMs remain accurate and up-to-date typically involves direct fine-tuning of the model. While effective, this method can be time-consuming and risk degrading performance on tasks the model has previously learned. To overcome these challenges, the concept of Model Refinement (MR) [151] has been developed as a technique to correct specific errors without disrupting the primary functionalities of the model. This approach, initially guided by a “reliability-locality-efficiency” principle, implements an efficient editing process using gradient descent. Subsequent research [152, 153] have expanded on this foundation, applying gradient decomposition techniques to larger models such as GPT-J-6B and T5-XXL. These methods focus on updating a subset of model parameters to modify the labels of specific inputs and integrate editing through retrieval mechanisms.
@@
Figure 17: An example of instruction following data.
Model Alignment. Model Alignment (MA) is a fundamental concept in the deployment of AI systems, aimed at ensuring that the actions and outputs of AI models are in harmony with human values, ethics, and preferences. This process involves the use of mathematical models, algorithmic adjustments, and iterative feedback to refine AI behaviours, enabling the model to adapt continuously to a stream of incoming data and remain up-to-date.
@@
IV-B2 Parameter Efficient Fine-tuning
@@
IV-B3 Other Fine-tuning
To address the growing computational and memory challenges of fine-tuning LLMs, recent research has proposed innovative approaches.
Density Efficient Fine-tuning. With increasing model sizes, traditional fine-tuning of all parameters becomes increasingly challenging. PEFT methods have gained popularity for effectively adapting pre-trained LMs. Recent research has identified the presence of activation sparsity within the Multilayer Perceptron (MLP) blocks of transformers. Lower activation density enables efficient model inference on hardware optimized for sparsity. Building on this insight, a new method called DEFT [166] has been proposed to reduce loss by promoting higher activation sparsity (or lower activation density) in pre-trained models.
Memory Efficient Fine-tuning. Traditional fine-tuning methods rely on backpropagation algorithms. However, as the model size increases, they demand exponentially more memory. Zeroth-order methods theoretically estimate gradients with only two forward passes but are typically very slow when optimizing LMs. MeZO [167] enhances the traditional Zeroth-order Stochastic Gradient Descent method by adapting it to in-place operations, maintaining the memory usage for fine-tuning LMs at the same level as that during inference.
IV-C LMs as the Non-Tuning Predictor
@@
Figure 20: In-context Learning for classification tasks.
IV-C1 Prompt-based Methods.
A notable method translates natural language instructions into executable robot actions [101]. This approach proposes customizable input prompts that seamlessly integrate into robotic execution systems or visual recognition programs, facilitating multi-step task plans while mitigating token limitation constraints of ChatGPT. In autonomous driving, LimSim++ [168] provides diverse types and modalities of prompt inputs, extracting road network and vehicle surroundings information at each decision frame, which is then conveyed in natural language to the driving agent (e.g., GPT-4).
IV-C2 Tokenizer-based Methods.
The introduction of LiDAR-LLM [169] leverages LLMs’ reasoning capabilities to comprehend outdoor 3D scenes fully. Addressing the scarcity of paired 3D LiDAR and text data, a three-stage training strategy generates datasets that align the 3D modality with the language embedding space of LLMs. Initially, LLM tokenizers were not designed to handle numerical data or recognize associations in time series effectively. This challenge was addressed through lightweight embedding layers and refined prompt engineering [170]. LLMTime [171] utilizes innovative tokenization techniques, transforming tokens into flexible, continuous-valued representations, enabling unadjusted LLMs to achieve or surpass traditional performance in zero-shot prediction tasks.
@@
Introduced in GPT-3 [18], ICL, initially popularized by GPT-3, enables LLMs to make accurate predictions using minimal demonstrations, bypassing the need for gradient updates. This update-free feature not only lowers adaptation costs but also broadens the model’s applicability across various tasks. The efficacy of ICL hinges on both the pre-training of the model on extensive datasets and the careful selection of demonstrations for task-specific predictions. Studies have identified several critical factors influencing the performance of ICL, including the label space, input text distribution, and the format of demonstrations [172]. Moreover, structural similarity, diversity, and complexity within demonstrations are also significant [173]. Recent research by Coda-Forno et al. [174] has shown that the capabilities of LLMs in ICL can be recursively enhanced through the ICL process itself. Fig. 20 shows an example of in-context learning. Several demonstrations containing both sources (comments for a book) and target classes (sentiments for the book) are fed into the LLM with the test input/query that only includes comment. The LLM can learn from the demonstrations and predicts the result without updating its parameters.
V IIoT: Connectivity Nexus for Emergence
With the ongoing development of the IIoT, its provision of stable and reliable connectivity has created new opportunities for resource-constrained devices to support multitasking [175]. Conventional IoT devices, due to their limited computational resources, are typically restricted to handling a single task at a time. For example, drones often face challenges in concurrently performing tasks such as vehicle tracking, traffic accident detection, and area intrusion monitoring. Recent advancements in modular deep learning, combined with the robust connectivity enabled by IIoT, have facilitated the execution of multiple tasks on a single resource-limited device, thus addressing longstanding limitations in multitask processing within such environments.
Figure 21: IIoT-enabled framework for modular routing and model merging on drones, enabling efficient multitasking.
V-A Modular LMs
In the IIoT context, the volume and variety of data generated by devices are enormous. Modular LMs enable specific modules to learn and optimize the processing of certain types of data. For example, one module may specialize in processing audio data from production lines, while another focuses on analyzing visual data. This specialization not only enhances data processing efficiency but also facilitates the reuse of learned knowledge across modules, reducing resource consumption and improving overall system performance.
V-A1 Modularization Facilitates Positive Transfer
By modularizing LMs, specific tasks or domain knowledge can be encapsulated within independent modules. This encapsulation not only preserves task-specific knowledge but also promotes positive transfer—where learning in one context enhances performance in another related but different context. This strategy effectively isolates and utilizes domain-specific insights, minimizing interference caused by irrelevant data and reducing the risk of negative transfer, where performance in one task could be adversely affected by learning from another.
V-A2 Achieving System Generalization
The modular architecture also supports system generalization by decoupling computation from routing. This allows the model to dynamically select and combine modules based on the incoming data. For instance, in the given scenarios, an LM on a drone might connect to different data providers or model providers through IIoT to enhance its capability on a single task. Moreover, this architecture enables the integration of functionalities to execute multiple tasks simultaneously using the same system infrastructure, such as a drone adapting its modules based on various sensor inputs and immediate demands, from monitoring crop health to anticipating industrial accidents.
Fig. 21 demonstrates how the integration of modular deep learning with the stable connectivity provided by the IIoT enables resource-constrained devices, such as drones, to execute multiple complex tasks simultaneously. Traditionally, IoT devices are limited by computational resources and can only handle single tasks, such as vehicle tracking or accident detection. However, leveraging IIoT’s reliable connectivity, data collected from diverse sensors (e.g., images, LiDAR, and maps) can be dynamically routed to task-specific expert models using modular routing mechanisms. Furthermore, model merging techniques consolidate the functionalities of multiple models into a single framework, reducing computational overhead while enhancing multitasking capabilities. Supported by IIoT’s distributed infrastructure, this framework facilitates efficient real-time execution of tasks, including surveillance, accident anticipation, and trajectory analysis, providing a scalable and intelligent foundation for industrial applications.
Research in this area primarily focuses on two key techniques: model routing and model merging.
V-B Model Routing in IIoT
Model routing refers to dynamically selecting the appropriate module or expert model based on the characteristics of the input data. This mechanism is analogous to the thalamus in the human brain, which routes sensory inputs to specific cortical regions. Routing offers the advantage of task-specific module selection, improving computational efficiency and reducing costs. Key advancements in this area include:
  • Spatiotemporal Routing: Temporal Netgrid models [176] optimize routing paths in satellite networks by predicting data transmission delays based on spatiotemporal characteristics.
  • To select the most suitable LLM for a new task from a pool of candidates, a router model has been proposed [177]. This data-driven approach leverages information from benchmark datasets to effectively guide the selection of the optimal LLM.
  • Embedding Space-Based Routing: Utilizing nearest-neighbor algorithms to predict model performance based on the distance between task samples in embedding space, enabling dynamic model selection [164].
  • Context-Aware Routing: The Tryage framework employs a routing mechanism that selects the best expert model based on user prompts and constraints such as model size and latency, achieving superior performance on heterogeneous datasets [178].
  • Hard and Soft Routing: Modular deep learning [179] leverages hard routing (binary selection) and soft routing (probabilistic selection) to optimize computational efficiency.
V-C Model Merging in IIoT
Model merging integrates the capabilities of multiple models into a single framework, reducing the computational and storage overhead required for multitasking. Key advancements in model merging include:
  • TIES-Merging: A weighted averaging strategy resolves interference among models to improve performance on tasks like text classification and question answering [180].
  • LoRM: A closed-form merging approach optimizes parameter-efficient modules (e.g., LoRA) through iterative freezing and merging cycles [181].
  • Dataless Fusion: This method averages model weights without additional training data, enhancing performance on downstream tasks [182].
  • Task Arithmetic: Task vectors are linearly combined to integrate multiple task functionalities into a single model, supporting task-specific updates and targeted adjustments [183].
  • Twin-Merging: This approach dynamically integrates expert models by learning twin vectors and routing networks, achieving superior performance on benchmark datasets [184].
Discussion. Model routing and model merging are complementary techniques that enable multitasking on resource-constrained devices. While routing dynamically assigns tasks to optimal modules, merging consolidates redundant models to minimize computational overhead. We summarize the work on model routing and model merging discussed in this paper in Table IX, providing a structured overview of these techniques. Future research can explore more efficient merging strategies and adaptive module integration in dynamic environments, further advancing intelligent applications.
TABLE IX: Modular LMs in IIoT.
Research Method Routing Merging Features
Tryage [178] $\checkmark$ Context-aware routing, inspired by the thalamus in the human brain.
Router [177] $\checkmark$ Leverage information from benchmark datasets to guide the selection of LLMs.
L2P [164] $\checkmark$ Routes based on the distance between new task samples and existing samples in the embedding space.
TNM [176] $\checkmark$ Used in satellite networks, selects routing paths based on spatiotemporal characteristics.
ModDL [179] $\checkmark$ Comprehensive discussion of hard and soft routing mechanisms, e.g., linear hypernetworks.
TIES-Merging [180] $\checkmark$ Resolves model merging conflicts, uses weighted averaging.
LoRM [181] $\checkmark$ Designed for parameter-efficient merging, e.g., LoRA modules.
RegMean [182] $\checkmark$ Merges model weights using weighted averaging without requiring additional training data.
Task Arithmetic [183] $\checkmark$ Integrates task-specific functionalities into a single model via task vector combinations.
Twin-Merging [184] $\checkmark$ Dynamically integrates expert models using learned twin vectors.
V-D Edge Test Time Adaptation
V-D1 Test Time Adaptation
Figure 22: Test Time Adaptation.
In Edge AI applications, model performance often deteriorates due to distribution shifts between training and testing data, which leads to suboptimal performance in the target domain. As shown in Fig. 22, Test Time Adaptation (TTA) has emerged as an effective solution [185]. TTA leverages test data to perform online adjustments of model parameters, aiming to enhance adaptability in dynamic environments. In the context of edge computing, the heterogeneous and dynamic nature of data distributions further amplifies this challenge, making TTA a critical research focus in the field.
Currently, research on TTA primarily focuses on several key aspects. First, to address the constraints of limited computational resources on edge devices, researchers aim to design efficient model update strategies. For instance, some studies propose updating only specific parameters [186, 187], such as those in normalization layers, to significantly reduce computational and memory overhead. Other approaches selectively upload valuable samples to the cloud for model updates, thereby reducing communication costs while improving update efficiency [188].
However, TTA processes may encounter catastrophic forgetting, where the model loses previously learned knowledge while adapting to new distributions. Continual Test-Time Adaptation addresses the challenge of adapting models to continuously evolving target domains. The introduction of the Visual Domain Adapter (ViDA) [189] allows the model to handle both domain-specific and shared knowledge. The Homeostatic Knowledge Allotment strategy optimizes the integration and utilization of knowledge from different ViDA, enhancing the model’s adaptability and preserving its plasticity across varying domains. Additionally, as TTA typically operates in an unsupervised manner, prediction confidence often fails to accurately reflect true model accuracy [190, 191]. To address this, some studies adopt model calibration techniques, such as consistency loss and min-max entropy regularization, to improve the alignment between predicted probabilities and actual performance.
Despite notable progress, several critical challenges remain in the field of TTA for Edge AI. Future research directions include developing lightweight TTA algorithms tailored for resource-constrained devices such as microcontrollers and sensor nodes; designing efficient TTA methods for specific applications, such as object detection, semantic segmentation, and speech recognition; and integrating TTA with federated learning frameworks to enable collaborative model updates across multiple edge devices, further improving generalization and robustness.
V-D2 Edge Cloud Collaboration
Edge-cloud collaboration addresses the challenges of deploying large models in resource-constrained edge environments by enabling cooperative inference and dynamic model adaptation. The Edge-Cloud Learning Model (ECLM) [192] employs a modular architecture, allowing personalized edge models to be derived from cloud-based large models. This approach minimizes parameter conflicts, providing efficient adaptation and scalability for diverse industrial applications.
Resource constraints often prevent the deployment of large models locally, resulting in inference delays [193] that hinder real-time IIoT applications such as autonomous driving, the metaverse [194], and digital twins [195]. To address this, edge-cloud collaborative methods leverage cooperative inference between edge and cloud models [196, 64]. By employing hard example mining, complex tasks are selectively offloaded to cloud-based large models for detailed processing, while simpler tasks are handled locally by lightweight edge models. This strategy ensures low latency for real-time applications while improving the accuracy of challenging tasks.
As illustrated in Fig. 23, cloud-based large models are critical in handling complex scenarios, such as high object density or adverse visual conditions in edge-cloud collaborative systems. By offloading computationally intensive tasks to the cloud, these systems enhance overall performance, improving accuracy, reliability, and safety.
Recent frameworks emphasize integrating resources to balance responsiveness and computational capability. The Cloud-Edge Collaborative Computing (CECC) framework [197] combines back-end and edge resources to deliver flexible and efficient services. For example, the EdgeFM system [198] uses cloud-based Feature Maps to enhance real-time model adaptation, maintaining high accuracy under dynamic conditions. Similarly, the EVOLVE framework [199] advances cloud-edge synergy through multiple cloud-based expert models that refine self-supervised learning on edge devices. Its dynamic expert aggregation mechanism adjusts model weights in real time based on reliability and expertise, effectively managing dynamic data streams and sustaining high performance in complex environments.
Figure 23: Edge-Cloud Collaborative Driving Planning with LMs.
VI IIoT: Evolution Pathway for LMs
In IIoT environments, the dynamic nature of data necessitates models that can continuously adapt to evolving scenarios by learning new knowledge. The relevance of continual learning refers to the interconnection between continual learning mechanisms and IIoT environments and applications. Traditional machine learning methods rely on retraining with large datasets, which is both costly and inefficient in IIoT contexts. This inefficiency arises due to the inability of traditional methods to address challenges such as data heterogeneity, resource constraints, and dynamic data streams. Continual learning overcomes these challenges by incrementally integrating new data into existing knowledge, offering a more efficient and cost-effective solution. Furthermore, continual learning enhances model adaptability and scalability. Its applications in IIoT are extensive, spanning areas such as predictive maintenance [14, 47], intelligent manufacturing, and smart transportation.
VI-A Theoretical Insights